CRY+VIS Interface Meeting
Date: 2021/11/12 13:00-13:50
Participants: Takahashi, Ikeda, Yasui, Sato, Kimura, Uchiyama, Tomaru, Aso, Sugiyama, Ushiba, Miyoki, Akutsu, Node
- IX: Health check. Height adjustment. Investigation of the F0 behavior.
- EX: Health check.
- IY: Cross (Flip) cables were replaced.
- EY: Preparation of ACCs.
- Replacement of Flip cable: IX/IY on 11/12.
- TM height setting: ITMX on 11/9.
- Oplev: Setup of the TM Oplev in IY after the height adjustment.
- ACC: Ready to install to EX. Measurement of ACCs for IY is on going. Actuator of H1 ACC was too weak. It should be modified.
- Commissioning in air: Health check. Review was held on 11/8.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Nov. 24-25 in IXV.
- Heater control: Power driver was installed to EY.
- Power source for Oplev: Implementing.
Date: 2021/11/12 14:00-14:37
Participants: Takahashi, Aso, Fabian, Sato, Miyoki, Ushiba, Sugiyama, Tanaka, Terrence
- Adjusted the height of TM in IX. The F0 LVDT was recalibrated.
- Investigating the coupling from the F0 actuator to the IP motion.
- Measured the TF of ACCs for IY. Actuator of H1 ACC was too weak. It should be modified.
- Aluminium windshields were built for SR3, SR2 and BS.
- Terrence did some measurements of the improvement produced by the windshield.
- Yasui-san and Terrence calibrated the BS oplev QPDs. They aimed to keep the BS pitch and yaw reference following a procedure.
- We began changing the flip adapters in SR3 today.
- On Monday we finished changing the flip adapters and measuring the resistances, etc. in PR2. Compilation of the data of the three Type-Bp suspensions is pending.
- On Monday and Tuesday we lowered the SF keystone.
- We installed new hardware to support ballast mass on top of the keystone.
- In total we added 399 grams. We were aiming for 405 grams.
- The temperature was very high and different to what it was when we did the initial assessment. Later, when we get the O3GK temperature back we will have to adjust the mass again.
- We lowered the yoke-coil assembly and the BF Damper by 5.2 mm also. We used the pitch of threads to do this and confirmed with the caliper in the case of the yoke-coil assembly.
- We used the SF fishing rod to set the mirror at the correct height. The same setposition happened to be a suitable position, yielding the mirror 0.2 mm above the expected beam height. This is not such a surprising result because we lowered the yoke-coil assembly 5.2 mm also.
- Unfortunately, during out work on Monday we disturbed the pitch of the mirror.
- On Wednesday we corrected the pitch.
- All evidence indicated this disturbance was similar to the previous ones: the mirror and IM moved in similar ways as before, but we didn't check the details.
- With the BF keystone at -300 um away from the setpoint, V1 OSEM readout was more than 11,000 counts, which is roughly what I was expecting. We had left it in 9,400 cnt. last time we adjusted it.
- After correcting the pitch and moving BF to the setpoint V1 OSEM readout became 1,300 cnt. It's not excellent but it's not bad!
- The transfer functions were fine, although one feature chaged in the vertical transfer functions IM-V, BF and SF: a feature dissapeard at 260 mHz and a peak appeared at 343 mHz. This is likely the effect of lowering the SF keystone.
- Other than this effect, the system looks healthy.
- They were able to engage the control system and align the green laser.
- Modified the PR3 damping control.
- Type-A team: Investigation of the coupling in IX. Preparation for the ACCs in EY.
- Type-B team: SR3 TM coil. BS LVDT. PR3 height.
(Date in Kamioka)
The week from 11/15
- Takahashi: 11/15-19
- Sato: 11/15-19
- Hirata: 11/16-19
The week from 11/22
- Takahashi: 11/24-26